Cos 315º = √2 (b) tanpa tan= k tansy = 30º = (v2 §) (c) sec (-120°)

evaluate the following without calculator. Give exact values.
Answer #1

tam sini = tan(1-1) and = -tan 176 [in quadrant ehen - Seto. © = - ta tanoz-ve] See (-120°) = See (120°) ; , Eisee (-0) = cos